OpenStudy (anonymous):

If f(x) = 2x^2 – 4x + 5 and g(x) = 2x – 2, what is f(x) + g(x) ? 2x^2 + 2x – 3 2x^2 + 6x – 7 2x^2 – 6x + 7 2x^2 – 2x + 3

OpenStudy (anonymous):

well if this helps. add -4x+5 to 2x-2 and the result will be the equation

OpenStudy (anonymous):

fx)+g(x)=2x^2-2x+3

OpenStudy (anonymous):

your only adding the like numbers between the two

OpenStudy (anonymous):

-4x+5 2x-2 you see how they both look similar. so you can add them. or "combine likes terms".

OpenStudy (anonymous):

ok hold on lemme get some paper

OpenStudy (anonymous):

2x^2 isnt addable because no other number has x^2. so you leave that alone

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[2x ^{2}-4x+ 5+2x-2\]

OpenStudy (anonymous):

add that together? ^^

OpenStudy (anonymous):

so when you write it out it all looks like one expression you see? so you just combine "like terms" and as you see 4x is negative but 2x is positive. 5 is positive and 2 is negative so you have to do the math.

OpenStudy (anonymous):

you write it out. and you just add the terms that look alike.

OpenStudy (anonymous):

thank you for helping me

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

no still writing one sec

OpenStudy (anonymous):

ok... 2x^2 – 2x + 3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

can u get this? 8x+4-2x-2

OpenStudy (anonymous):

lemme try

OpenStudy (anonymous):

when i combine 8x and 2x do i include the negative sign or no?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

k

OpenStudy (anonymous):

idk wth i just did XD

OpenStudy (anonymous):

6 + 6 can't be the answer

OpenStudy (anonymous):

8x + -2x = 6x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and then 4 + 2 = 6

OpenStudy (anonymous):

i dunno im an idiot XD sorry

OpenStudy (anonymous):

4-2=2 if there is a negative sighn in front of the 2 then it is negative. just how 2x was negative

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

8x+4-2x-2

OpenStudy (anonymous):

8x-2x 4-2

OpenStudy (anonymous):

so its...6x + 2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

:D YUSH!!

OpenStudy (anonymous):

thanks dude :P

OpenStudy (anonymous):

no problem